OpenStudy (anonymous):

(7x-2)(3x^2-4x-6) Would someone kindly help me to compare answers? Thanks in advance.

OpenStudy (anonymous):

depends on what u r trying to do

OpenStudy (anonymous):

multiplying polynomials

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

Distribute!

OpenStudy (anonymous):

yes pol

OpenStudy (anonymous):

i would probably seperate it into: 7x(3x^2-4x-6)-2(3x^2-4x-6) and go from there

OpenStudy (anonymous):

\[(7x-2)(3x^2 - 4x - 6)\]\[=3x^2(7x-2) - 4x(7x-2) - 6(7x-2)\]\[=...\]

myininaya (myininaya):

\[(ax^2+bx+c)(dx^2+ex+f)=ax^2(dx^2+ex+f)+bx(dx^2+ex+f)+c(dx^2+ex+f)\]

OpenStudy (anonymous):

Either way

OpenStudy (anonymous):

i got 21x^3-34X^2+34x+12

myininaya (myininaya):

\[=ax^2(dx^2)+ax^2(ex)+ax^2(f)+bx(dx^2)+bx(ex)+bx(f)+c(dx^2)+c(ex)+c(f)\]

OpenStudy (anonymous):

you are correct.

OpenStudy (anonymous):

thanks!

myininaya (myininaya):

anyways we can do \[7x(3x^2-4x-6)-2(3x^2-4x-6)\] or \[3x^2(7x-2)-4x(7x-2)-6(7x-2)\]

OpenStudy (anonymous):

I disagree with your x term.

OpenStudy (anonymous):

i used FOIL

OpenStudy (anonymous):

Oh wait, no I don't

OpenStudy (anonymous):

Oh wait, yes I do.

OpenStudy (anonymous):

21x^3-34x^2-34x+12

OpenStudy (anonymous):

yeah i missed the minus sign

myininaya (myininaya):

\[=7x(3x^2)-7x(4x)-7x(6)-2(3x^2)-2(-4x)-2(-6)\] \[=21x^3-28x^2-42x-6x^2+8x+12=21x^3-34x^2-34x+12\]